OpenStudy (anonymous):

let -pi/2 < x < pi/2 f(x)=e^x*sin(x) (a) Find the gradient of y=f(x) when x=pi/4 (b) Find the coordinates of the point where the gradient is zero I know part (a) but im struggling in part (b). Could someone please help explain it to me. Thx

OpenStudy (loser66):

Is it not that gradient mean derivative??

OpenStudy (loser66):

and for part b, just let it =0 solve for x then plug back to y to solve for y.

OpenStudy (loser66):

\[f'(x) = e^x(sinx + cosx)=0\]

OpenStudy (loser66):

\(e^x\neq 0~~\forall x\) --> sinx + cos x =0 only. I think you can step up from here, right?

OpenStudy (anonymous):

no thats exactly where i stopped :)

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

what do you mean? that is part b

OpenStudy (anonymous):

yes

OpenStudy (loser66):

you take gradient of f(x) means derivative of f(x) , right?

OpenStudy (anonymous):

its asking me to find the coordinates

OpenStudy (loser66):

yes, it is.

OpenStudy (anonymous):

the answer in the text book is (-pi/4, -sqrt2/2*e^-pi/4)

OpenStudy (loser66):

now solve for sinx + cos x=0 , you have x = -pi/4 for example

OpenStudy (loser66):

now, plug x = -pi/4 into f(x) you have f(x) = e^x sin x

OpenStudy (anonymous):

where did x=-pi/4 come from?

OpenStudy (loser66):

from sin x = -cosx

OpenStudy (anonymous):

could you please explain it to me step by step

OpenStudy (loser66):

ok, restart

OpenStudy (loser66):

what is the gradient of f(x)? in general??

OpenStudy (anonymous):

f'(x)

OpenStudy (loser66):

=?

OpenStudy (anonymous):

f'(x)=e^x(sinx+cosx)

OpenStudy (loser66):

yes, now, they ask " where it is =0 " means f'(x) =0 at x = somewhere, right? so, just let it =0 and solve for x

OpenStudy (anonymous):

e^x(sinx+cosx)=0

OpenStudy (loser66):

and e^x \(\neq 0\) for all x, right?

OpenStudy (anonymous):

right

OpenStudy (loser66):

so, to get it is =0 , you have only 1 option, that is sinx + cos x =0, right?

OpenStudy (anonymous):

sinx= -cos x

OpenStudy (loser66):

in the interval \([-\pi/2,\pi/2]\) , which value gives you sinx =-cos x??

OpenStudy (anonymous):

7pi/4

OpenStudy (loser66):

is it not that 7pi/4 = -pi/4??

OpenStudy (anonymous):

correct

OpenStudy (loser66):

or you can see it from  You have to take this value because 7pi/4 is out of the interval

OpenStudy (loser66):

got me so far?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

now, at x = -pi/4 , f' (x) =0 ( remember, just f' not f) so, if x = -pi/4, y =? ( y is f(x)

OpenStudy (anonymous):

do i subtract 2pi from it?

OpenStudy (loser66):

no need to do, just take the logic of 7pi/4 = -pi/4

OpenStudy (anonymous):

ok

OpenStudy (loser66):

ok, answer me, x = -pi/4, f(x) =?

OpenStudy (anonymous):

i plug it in

OpenStudy (loser66):

yup

OpenStudy (anonymous):

i get -sqrt2/2*e^-pi/4

OpenStudy (anonymous):

thats the exact answer in the textbook

OpenStudy (anonymous):

thank you so much for your help and time

OpenStudy (loser66):

You got the concept, right?

OpenStudy (anonymous):

yes :)

OpenStudy (loser66):

good